A 73 kg object slowed down to 3.0 m/s. If the work done on this object is -2.0 x 10^3 J, then what is the initial speed
1
Expert's answer
2020-05-11T20:20:38-0400

According to the kinetic energy theorem, the work was spent to slow the body down:


"KE_f-KE_i=W,\\
\frac{1}{2}mu^2-\frac{1}{2}mv^2=W,\\
\space\\
\frac{1}{2}mv^2=\frac{1}{2}mu^2-W,\\
\space\\
v=\sqrt{u^2-\frac{2W}{m}}=\sqrt{3^2-\frac{2(-2.0 \cdot 10^3)}{73}}=8\text{ m/s}."